A firm is must choose to buy the GSU-3300 or the UGA-3000. Both machines make the firm’s production process more efficient which in turn increases incremental cash flows. The GSU-3300 produces incremental cash flows of $26,871.00 per year for 8 years and costs $99,126.00. The UGA-3000 produces incremental cash flows of $28,696.00 per year for 9 years and cost $125,695.00. The firm’s WACC is 8.83%. What is the equivalent annual annuity of the GSU-3300?
Answer #1

The first step is to calculate the present value of cash flows of GSU-3300:

PV of cash flows = Incremental cash flows * (1 - (1+R)^(-n)) / R

where R = 8.83% and N = 8

= 26871 * (1 - (1+8.83%)^-8) / 0.0883

= 149670.47

NPV = PV of cash flows - Initial investment = 149670.47 - 99126

= 50544.47

Equivalent annual annuity = NPV * R / (1 - (1+R)^(-n))

= 50544.47 * 8.83% / (1 - (1+8.83%)^(-8))

= 9074.47